OpenStudy (anonymous):

Trying to classify the roots as (a) 2 real(uneaqual roots), (b) 1 double root or (c) 2 imaginary conjugate roots x^2+4x=5

OpenStudy (anonymous):

x^2+4x=5 x^2+4x-5 =0 x^2+(-x+5x)-5 =0 x(x-1)+5(x-1)=0 (x-1)(x+5) =0 x=1 or x=-5 2 real unequal roots

OpenStudy (anonymous):

thank you